The formula b(t) = 1.353(1.9025)^t represents the number of broadband internet users (in millions) where t is the time in years measured from 1998. How many broadband users (to the nearest million) does this formula predict for 2004?
Answer by checkley77(12844)

b(t) = 1.353(1.9025)^t
t=2004-1998=6
b(t)=1.353(1.9025)^6
b(t)=1.353*47.4185
b(t)=64 million broadband users in 2004 is the projection.